# A Modern Look at the Oscillation Physics Case for a Neutrino Factory

### v2

@@ 3.2 Sensitivity to other oscillation parameters
In fig. 6 we show the precision on $\sin^{2}\theta_{23},~{}\sin^{2}2\theta_{13}$ and $\Delta m_{31}^{2}$ [^7] Just like DUNE and HK a NF has only very limited sensitivites to the solar parameters Denton:2023zwa . assuming 10 years of a NF with $10^{21}$ muon decays per year combined with 10 years of DUNE and HK each. We compare the results to existing constraints and the expected sensitivies from DUNE and HK alone, and 20 years of DUNE and 10 years of HK. We summarize the results in tab. 2. In fig. 7 we show the correlations between $\sin^{2}\theta_{23}-\Delta m_{31}^{2}$ and $\theta_{13}-\delta$.
We find that the combination of DUNE and HK will drastically improve the precision on the atmospheric parameter $\Delta m_{31}^{2}$ and $\theta_{23}$. The addition of a NF will further improve the precision nearly independent on the assumption of CID. The improvement on the precision on $\Delta m_{31}^{2}$ is slightly larger for the BNL setup where it can also improve over 20 years of DUNE combined with 10 years of HK. For $\theta_{23}$ the improvements in precision due to the addition of a NF are most noticeable for $\sin^{2}\theta_{23}^{\text{true}}=0.5$ where a NF leads to a more symmetric precision. Again we find that the BNL setup leads to slightly higher precision that the FNAL setup and can even improve over 20 years of DUNE plus 10 years of HK. Finally, DUNE and HK lead to slight improvement on the precision of $\sin^{2}2\theta_{13}$ which can even be further increased by the addition of a NF. Both setups could even improve over 20 years of DUNE and 10 years of HK with additional improvements coming from CID of either electrons or muons. Also the two-dimensional precision on $\Delta m_{31}^{2}-\sin^{2}\theta_{23}$ and $\theta_{13}-\delta$ benefit from the addition of a NF to DUNE and HK.
@@
Figure 6: Precision on $\Delta m_{31}^{2},~{}\sin^{2}(2\theta_{13}),~{}\sin^{2}\theta_{23}$ for true NO. For $\theta_{23}$ we assume three different true values $\theta_{23}^{\text{true}}=(42^{\circ},~{}45^{\circ},~{}49^{\circ})$. We show the results for 10 years of NF with $10^{21}$ muon decays per year and a 40 kT FD and $E_{\mu}=5~{}(8)$ GeV for the FNAL (BNL) setup on the left (right) plots, and for different assumptions on CID. The results for the NF also include 10 years of both HK and DUNE. We also compare the results to 20 years of DUNE in addition to 10 years of HK. The gray curves show the current precision on these parameters from Esteban:2020cvm ; DayaBay:2022orm .
Figure 7: Two-dimensional precision on the oscillation parameters assuming DUNE+HK+NF (10 years of each) with the FNAL-SURF baseline assuming different true values marked with a star. The contours are drawn at 90% C.L. for 2 dofs.
4 Discussion and Recommendations
@@
We now highlight the differences between a neutrino factory and DUNE as presented here. First, the shape of the spectrum is quite different and it tends to rise to the high energy part of the spectrum. This is compared to the broader spectrum that DUNE will see based on the specific horn configuration designed to have good sensitivity to $\delta$ DUNE:2020ypp ; Calviani:2014cxa . Second, the spectrum of neutrinos for a neutrino factory is precisely predicted, while that for DUNE carries significant uncertainties. Third, a neutrino factory provides access to more channels (notably $\nu_{e}\to\nu_{e}$, $\nu_{e}\to\nu_{\mu}$, and their conjugates) than are regularly easily accessible. Fourth, for a given accelerator complex, the flux of neutrinos from a neutrino factory will be similar to, but likely less, than that for DUNE. The exact details here depend on muon cooling efficiencies and the geometry of the neutrino factory, but all factors in this point are no more than $\mathcal{O}(1)$ factors.
Our results show that while CID does enhance the oscillation parameters’ precision, CID plays a smaller role than emphasized in studies a decade ago. This is primarily due to the fact that LArTPC detectors have sufficiently good energy resolution to distinguish signal from background as well as identify the different flavors. In fact, the LArTPC energy resolution could be even better than assumed here, see e.g. DeRomeri:2016qwo ; Friedland:2018vry ; Kopp:2024lch .
@@
Using a similar detector setup for a neutrino factory as for near-future experiments could further reduce the uncertainties on the cross section and detector efficiencies, as there is an ongoing program to measure the neutrino cross section on Ar and advancing the LAr technology Machado:2019oxb .
A neutrino factory could be an appealing possible future neutrino oscillation experiment should the results of HK and DUNE disagree. If the weak tension between NOvA and T2K in the CP violation measurements Denton:2020uda ; Chatterjee:2020kkm persists, DUNE will be able to probe it Denton:2022pxt . However if the tension requires further oscillation studies at a higher neutrino energy, longer baseline, and overall smaller flux uncertainties than fixed target oscillation experiments a neutrino factory is a favorable setup for future studies. [^8] Note that atmospheric neutrino experiments can also provide oscillation studies at higher energies and longer baselines however their flux uncertainty and flavor composition is more uncertain than at a neutrino factory. Furthermore, at a neutrino factory the neutrino energy is flexible and tunable Delahaye:2013jla . Finally, at a neutrino factory six oscillation channels and their CP conjugate ones are accessible with similarly large number of events as the initial neutrino beam consists equally of $\nu_{e},\nu_{\mu}$. The presence of numerous oscillation channels provides multiple independent means of getting at the oscillation parameters, each with different dependences on the systematic uncertainties.
Table 1: Resolution on $\delta$ for combinations of different experiments for two true values of $\delta$. For DUNE we assume 480 kT-MW-yr, for HK we assume 2.47 MT-MW-yr, for the neutrino factory exposure we assume a 40 kT LAr detector and a total of $10^{22}$ muon decays – nominally 10 years of each experiment unless otherwise specified. We also show the sensitivities for 960 kT-MW-yr of DUNE, twice its nominal exposure, combined with 2.47 MT-MW-yr of HK.
@@
Table 2: Resolution on the other parameters for combinations of different experiments. For DUNE we assume 480 kT-MW-yr, for HK we assume 2.47 MT-MW-yr, for the neutrino factory exposure we assume a 40 kT LAr detector and a total of $10^{22}$ muon decays – nominally 10 years of each experiment unless otherwise specified. We also show the sensitivities for 960 kT-MW-yr of DUNE, twice its nominal exposure, combined with 2.47 MT-MW-yr of HK.
@@ 5 Conclusions
Motivated by the recent P5 report, we study the physics potential of a neutrino factory in improving the precision on the determination of the remaining undetermined neutrino oscillation parameters with a focus on the complex phase $\delta$ which governs the amount of CP violation in the neutrino sector. We consider two different baselines and muon energies with a LAr detector based at SURF. Similar baselines and energies were considered as a low-energy neutrino factory Geer:2007kn ; Bross:2007ts appropriate to measure CPV for large values of $\theta_{13}$ Tang:2009wp ; Dighe:2011pa ; Ballett:2012rz ; Christensen:2013va . Such a neutrino factory setup could be achieved with a proton beam with 2-4 MW Holtkamp:2000xn ; Ozaki:2001bb ; NeutrinoFactory:2004odt ; Geer:2006tb ; Ankenbrandt:2009zza .
@@ Appendix B Results for inverted ordering and higher energies for BNL setup
Table 3: Results for precision on parameters for BNL baseline at higher energies. For DUNE we assume 480 kT-MW-yr, for HK we assume 2.47 MT-MW-yr, for the neutrino factory exposure we assume a 40 kT LAr detector and a total of $10^{22}$ muon decays, nominally 10 years of each experiment.

### v3

@@ 3.2 Sensitivity to other oscillation parameters
In fig. 6 we show the precision on $\sin^{2}\theta_{23},~{}\sin^{2}2\theta_{13}$ and $\Delta m_{31}^{2}$ [^8] Just like DUNE and HK a NF has only very limited sensitivites to the solar parameters [10]. assuming 10 years of a NF with $10^{21}$ muon decays per year combined with 10 years of DUNE and HK each. We compare the results to existing constraints from global analyses and Daya Bay, and the expected sensitivies from DUNE and HK alone, and 20 years of DUNE and 10 years of HK. We summarize the results in tab. 2. In fig. 7 we show the correlations between $\sin^{2}\theta_{23}-\Delta m_{31}^{2}$ and $\theta_{13}-\delta$.
We find that the combination of DUNE and HK will drastically improve the precision on the atmospheric parameter $\Delta m_{31}^{2}$ and $\theta_{23}$. The addition of a NF will further improve the precision nearly independent on the assumption of CID. We focus on cases where the octant of $\theta_{23}$ can easily be identified, but a NF will also increase the sensitivity to determining to correct octant for regions of parameter space close to maximal mixing. The improvement on the precision on $\Delta m_{31}^{2}$ is slightly larger for the BNL setup where it can also improve over 20 years of DUNE combined with 10 years of HK. For $\theta_{23}$ the improvements in precision due to the addition of a NF are most noticeable for $\sin^{2}\theta_{23}^{\text{true}}=0.5$ where a NF leads to a more symmetric precision. Again we find that the BNL setup leads to slightly higher precision that the FNAL setup and can even improve over 20 years of DUNE plus 10 years of HK. Finally, DUNE and HK lead to slight improvement on the precision of $\sin^{2}2\theta_{13}$ over Daya Bay which can even be further increased by the addition of a NF. Both setups could even improve over 20 years of DUNE and 10 years of HK with additional improvements coming from CID of either electrons or muons. Also the two-dimensional precision on $\Delta m_{31}^{2}-\sin^{2}\theta_{23}$ and $\theta_{13}-\delta$ benefit from the addition of a NF to DUNE and HK.
@@
Figure 6: Precision on $\Delta m_{31}^{2},~{}\sin^{2}(2\theta_{13}),~{}\sin^{2}\theta_{23}$ for true NO. For $\theta_{23}$ we assume three different true values $\theta_{23}^{\text{true}}=(42^{\circ},~{}45^{\circ},~{}49^{\circ})$. We show the results for 10 years of NF with $10^{21}$ muon decays per year and a 40 kT FD and $E_{\mu}=5~{}(8)$ GeV for the FNAL (BNL) setup on the left (right) plots, and for different assumptions on CID. The results for the NF also include 10 years of both HK and DUNE. We also compare the results to 20 years of DUNE in addition to 10 years of HK. The gray curves show the current precision on these parameters from [72, 73].
Figure 7: Two-dimensional precision on the oscillation parameters assuming DUNE+HK+NF (10 years of each) with the FNAL-SURF baseline assuming different true values marked with a star. The contours are drawn at 90% C.L. for 2 dofs. We assume true NO.
4 Discussion
@@
We now highlight the differences between a neutrino factory and DUNE as presented here. First, the shape of the spectrum is quite different and it tends to rise to the high energy part of the spectrum. This is compared to the broader spectrum that DUNE will see based on the specific horn configuration designed to have good sensitivity to $\delta$ [2, 74]. Second, the spectrum of neutrinos for a neutrino factory is precisely predicted, while that for DUNE carries significant uncertainties. Third, a neutrino factory provides access to more channels (notably $\nu_{e}\to\nu_{e}$, $\nu_{e}\to\nu_{\mu}$, and their conjugates) than are regularly easily accessible. Fourth, for a given accelerator complex, the flux of neutrinos from a neutrino factory will be similar to, but likely less, than that for DUNE. The exact details here depend on muon cooling efficiencies and the geometry of the neutrino factory, but all factors in this point are no more than $\mathcal{O}(1)$ factors.
Our results show that while CID does enhance the oscillation parameters’ precision, CID plays a smaller role than emphasized in studies a decade ago. This is primarily due to the fact that LArTPC detectors have sufficiently good energy resolution to distinguish signal from background as well as identify the different flavors. In fact, the LArTPC energy resolution could be even better than assumed here, see e.g. [75, 76, 77].
@@
Using a similar detector setup for a neutrino factory as for near-future experiments could further reduce the uncertainties on the cross section and detector efficiencies, as there is an ongoing program to measure the neutrino cross section on Ar and advancing the LAr technology [78].
A neutrino factory could be an appealing possible future neutrino oscillation experiment should the results of HK and DUNE disagree. If the weak tension between NOvA and T2K in the CP violation measurements [79, 80] persists, DUNE will be able to probe it [81]. However if the tension requires further oscillation studies at a higher neutrino energy, longer baseline, and overall smaller flux uncertainties than fixed target oscillation experiments a neutrino factory is a favorable setup for future studies. [^9] Note that atmospheric neutrino experiments can also provide oscillation studies at higher energies and longer baselines however their flux uncertainty and flavor composition is more uncertain than at a neutrino factory. Furthermore, at a neutrino factory the neutrino energy is flexible and tunable [82]. Finally, at a neutrino factory six oscillation channels and their CP conjugate ones are accessible with similarly large number of events as the initial neutrino beam consists equally of $\nu_{e},\nu_{\mu}$. The presence of numerous oscillation channels provides multiple independent means of getting at the oscillation parameters, each with different dependences on the systematic uncertainties.
Table 1: Resolution on $\delta$ for combinations of different experiments for two true values of $\delta$ in NO. For DUNE we assume 480 kT-MW-yr, for HK we assume 2.47 MT-MW-yr, for the neutrino factory exposure we assume a 40 kT LAr detector and a total of $10^{22}$ muon decays – nominally 10 years of each experiment unless otherwise specified. We also show the sensitivities for 960 kT-MW-yr of DUNE, twice its nominal exposure, combined with 2.47 MT-MW-yr of HK.
@@
Table 2: The $1\sigma$ resolution on the other parameters for combinations of different experiments. For DUNE we assume 480 kT-MW-yr, for HK we assume 2.47 MT-MW-yr, for the neutrino factory exposure we assume a 40 kT LAr detector and a total of $10^{22}$ muon decays – nominally 10 years of each experiment unless otherwise specified. We also show the sensitivities for 960 kT-MW-yr of DUNE, twice its nominal exposure, combined with 2.47 MT-MW-yr of HK. The current uncertainties are $\sim 1^{\circ}$, $\sim 0.1^{\circ}$, and $\sim 20\times 10^{-6}$ eV2 for $\theta_{23}$, $\theta_{13}$, and $|\Delta m^{2}_{31}|$, respectively. We assume true NO.
@@ 5 Conclusions
Motivated by the recent P5 report, we study the physics potential of a neutrino factory in improving the precision on the determination of the remaining undetermined neutrino oscillation parameters with a focus on the complex phase $\delta$ which governs the amount of CP violation in the neutrino sector. We consider two different baselines and muon energies with a LAr detector based at SURF. Similar baselines and energies were considered as a low-energy neutrino factory [83, 45] appropriate to measure CPV for large values of $\theta_{13}$ [84, 85, 86, 87]. Such a neutrino factory setup could be achieved with a proton beam with 2-4 MW [88, 89, 90, 91, 92].
@@ Appendix B Results for inverted ordering and higher energies for BNL setup
In figs. 9, 10, 11 we show the corresponding results in IO to figs. 4, 6, 7 in the main text. The results are quite similar to the NO.
Table 3: Results for precision on parameters for BNL baseline at higher energies and true NO. For DUNE we assume 480 kT-MW-yr, for HK we assume 2.47 MT-MW-yr, for the neutrino factory exposure we assume a 40 kT LAr detector and a total of $10^{22}$ muon decays, nominally 10 years of each experiment.
@@
Figure 9: Precision on $\delta$ in degrees as a function of neutrino factory exposure and CID in IO. We assume true IO. The neutrino factory has a baseline and muon energy of 1284.9 km and 5 GeV and also includes 10 years of both DUNE and HK. The shaded regions are for electron CID and the lines are for muon CID; electron CID does slightly better in most cases. The horizontal lines show possibly achievable muon CID. The true values of $\delta$ are 00 (left) and $-90^{\circ}$ (right). We assume as ratio of neutrino to anti-neutrino running time of 1:1 for the neutrino factory and DUNE 1:1 and 1:3 for HK.
Figure 10: Precision on $\Delta m_{31}^{2},~{}\sin^{2}(2\theta_{13}),~{}\sin^{2}\theta_{23}$ for true IO. For $\theta_{23}$ we assume three different true values $\theta_{23}^{\text{true}}=(42^{\circ},~{}45^{\circ},~{}49^{\circ})$. We show the results for 10 years of NF with $10^{21}$ muon decays per year and a 40 kT FD and $E_{\mu}=5~{}(8)$ GeV for the FNAL (BNL) setup on the left (right) plots, and for different assumptions on CID. The results for the NF also include 10 years of both HK and DUNE. We also compare the results to 20 years of DUNE in addition to 10 years of HK. The gray curves show the current precision on these parameters from [72, 73].
Figure 11: Two-dimensional precision on the oscillation parameters assuming DUNE+HK+NF (10 years of each) with the FNAL-SURF baseline assuming different true values marked with a star. The contours are drawn at 90% C.L. for 2 dofs. We assume true IO.